• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 09:52:29 on November18,2002

    Entry number 90237

    keyword=Fac. Dev. update

    No beam since Sat. midnight, except for a shot of 10 minutes last night.
    The wire-chamber tests were abandoned, Bigbite removed, the Left HRS
    polarity is being switched, the target is being warmed up with scopes.
    We plan to lock up around noon. Then we will do accelerator tests until
    18:00, and finally SRC tests.